# pnasresearcharticle

### v2

@@ Abstract
In Fig. 4, we explore the immediate impact of a singular perturbation (a “kick”) to the system (52). The perturbation is implemented as either a discrete change in the infected population (Fig. 4A-C) or the susceptible population (Fig. 4D). Since only a single “kick” is given, the system will only transiently settle into a multiannual cycle and we thus measure the strength of the 5-year cycle during the 20 years immediately following the perturbation. Fig. 4A shows that the timing of the perturbation matters, with a kick during the low-transmission part of the season most likely to excite the multi-year cycle. In other words, a kick of a certain size (whether in the positive or negative direction) has a larger effect if it occurs when circulation is already low. However, Fig. 4B reveals that this seasonal dependence is completely accounted for by the relative magnitude of the perturbation, i.e. the change in incidence divided by the current incidence at that time. If plotted as a function of this relative magnitude, $\Delta I/I$, the strength of the excited cycle shows only a very faint seasonal dependence.
The annual as well as the (transient) five-year cycles are shown in the $(S,I)$ configuration space (with $R=1-S-I$ suppressed) in Fig. 4C. The rate at which $I(t)$ and $S(t)$ changes can be judged by the spacing between the colored markers, and the color of the markers themselves indicate the time of year. As is evident, the five-year cycle can be quite complex. The perturbation that initiated the five-year cycle is visible as the approximately vertical line connecting the annual and multiannual cycles.
Figure 3: Moderate environmental noise stimulates multiannual cycles, especially approximately 5-year cycles. A) Average Fourier spectrum across 1560 realizations (for each noise level indicated in the legend) of the prevalence $I(t)$ of the seasonal SIRS model with parameters determined by a fit to 2010-2025 data from Denmark. The noise source is random fluctuations in the transmission rate $\beta(t)$, implemented as Gamma-distributed multiplicative noise with mean $1$ and variance $\sigma^{2}$ (see Methods and (51)) The noise level shown is the value of $\sigma$. Each analyzed simulation timeseries was 100 years, and a 1000 year “transient” was discarded. B) Fourier spectra of four individual realizations at low to moderate noise levels.
Figure 4: One-off exogeneous perturbations (“kicks”) transiently excite multiannual cycles. A) After a discrete change in prevalence of magnitude $\Delta I$, the system may temporarily settle into a 5-year cycle. The 5-year cycle is particularly prominent if the perturbation happens near the time of lowest prevalence. The color indicates the power of the 5-year cycle in the Fourier spectrum, relative to the highest peak. The spectrum is based on the first 20 years post-perturbation. The red cross indicates the perturbation implemented in panel D. B) If the magnitude of the perturbation $\Delta I$ is proportional to the current prevalence $I(t)$, the prominence of the 5-year cycle ceases to depend on time of year. The red cross again indicates the perturbation implemented in panel D. C) Annual and 5-year cycles depicted in $(S,I)$ configuration space. The marker color denotes the time of year, and markers are one week apart. The approximately vertical line starting from the red cross shows the exogenous “kick” which causes a transition from an annual (thick yellow line) to a 5-year cycle (thin purple line). D) Perturbations to $S(t)$, rather than the prevalence.
@@
Although the deterministic seasonal SIRS model captures the 2010-2025 dynamics well, predicting or forecasting M. pneumoniae incidence may be challenging. We know that long-term predictions using the determinstic model are limited by its tendency to revert to annual dynamics. In the short term, however, a deterministic model may still be appropriate for forecasting. Fig. 5A shows the trajectory obtained by continuing the seasonal SIRS model for another 10 years (with parameters from the fit to 2010-2025 data). Supplementary Fig. S1 shows that the system in fact exhibits transient chaos, with a maximal local/instantaneous Lyapunov exponent (LLE) (53) that is mostly positive. A Lyapunov exponent $\lambda$ measures sensitivity to initial conditions – this also includes the tendency of the system to amplify ($\lambda>0$) or buffer ($\lambda<0$) perturbations from e.g. noise (54). The arithmetic mean value of the greatest LLE during the first two years is approx. $0.02/\text{wk}$. In practical terms, a separation between two trajectories will have grown by almost a factor of $10$ after the first two years of deterministic dynamics. LLEs were computed using the QR method described in (55).
@@
Moving beyond deterministic dynamics, the effects of environmental stochasticity on predictions are shown in 5B. Here, we assume a moderately high level of environmental stochasticity ($\sigma=0.15$, compare with the amplitude of seasonality at $0.19$) and explore the range of predicted outcomes. The fit to empirical data is assumed known, and noise is only introduced for the prediction part of the simulations. While sensitivity to stochastic effects may complicate medium-term predictions (note the spread between realizations in 5B from 2027 onward), some short-term features are relatively robust even at this noise level. Predictions agree that 2025 is likely to be a moderate-to-high incidence year, exemplifying how deterministic models yield valuable short-term forecasts even in transiently chaotic epidemiological systems. While long-term predictions using the deterministic model are limited by its decay into annual dynamics, the presence of stochasticity limits extended prediction not only by causing variability in the incidence during any given year, but also by enabling mode-hopping, i.e. transitions between annual and multiannual cycles (56, 49), as exemplified in Supplementary Fig. S6. That the predictions for the years immediately following the large NPI-induced perturbation agree well between the stochastic and deterministic simulations fits well with the general observation that restorative forces are stronger far from equilibrium leading the stochastic model to (temporarily) behave more like the deterministic model (42).
Figure 5: Prediction is hard. A) Prediction using deterministic model, parameters fitted to 2010-2025 data. B) Range of predictions given identical assumptions, but including noise in transmission ($\sigma_{\beta}=0.15$, 100 realizations). Noise can significantly affect long- and medium-term predictions, but all realizations agree that 2025 will see significant M. pneumoniae activity.
@@
In fitting the 2010-2025 time series, we used the PCR positivity rate as an incidence proxy, while we used the number of positive tests for the 1958-1995 data (total number of tests were not on record). Neither of these measures are perfect incidence proxies, and they suffer in distinct ways from biases introduced by varying testing intensities. If testing intensity is increased, this leads to an increase in the number of positive tests, while it leads to a decrease in the test positivity rate (since the denominator – the total number of tests – is increased). It may be possible to enhance fitting further by accounting for fluctuations in testing intensity when such data is available, something we have not attempted to do, beyond simple variance stabilization. Furthermore, the fidelity of the 1958-1995 data series is affected by both the (somewhat imperfect) test characteristics of the cold agglutination and complement fixation tests (since the samples considered positive are those which were positive in both of these tests (2)). We have not taken age structure into account (57), nor have we considered spatial effects which might enhance the impact of demographic stochasticity (58).
@@
The origins of non-trivial cycles in ecological systems and the extent to which stochasticity plays a role in shaping them are long-standing questions in ecology, including epidemiology. In (59), the authors used wavelet analysis to identify 3-4 year cycles in pertussis across 12 countries and noted that the periodicity was transient in nature, in line with results of (60, 61). While the origins of the multiannual cycles of M. pneumoniae are generally regarded as unclear (62, 21), a few different mechanisms of varying complexity have been presented in the literature. Zhang et al. (63) suggested a combination of strain dynamics, contact network structure and waning immunity as the explanation, while Omori et al. (64) explored three different possible explanations: school-term forcing, a low-dispersion distribution of immunity duration, and interference between M. pneumoniae serotypes. They find that, of these three explanations, a less variable duration of immunity (compared with the typically assumed exponential waning) best captures the generally observed patterns in that it allows for incidence oscillations.
Recently, the authors of (26) used a TSIR model to describe the post-COVID rebound of M. pneumoniae in Denmark as well as four UN regions (Europe, Asia, the Americas, and Oceania) but found that they required at least a 90% reduction in the transmission rate due to non-pharmaceutical interventions in order to fit the data.
@@
We find that intrinsic, noisy dynamics is sufficient to explain the complex cycles, offering a parsimonious model for the dynamics of M. pneumoniae. Earn et al. (65) demonstrated the existence of a rich landscape of stable cycles in the seasonally driven SEIR model at high transmission rates (basic reproductive numbers in excess of approx. 4). In the current case of M. pneumoniae however, multiannual cycles are unstable and require stochasticity or perturbations to be sustained and appear to be a result of stochastic resonance.
With the COVID-19 NPIs phased out, another major – if less abrupt – change in M. pneumoniae epidemiology is on the horizon. Macrolide resistance has been growing in recent years, albeit with large geographical differences (66), and it is currently unclear how this will affect epidemiological dynamics. More generally, there is a need for comparative studies of the factors governing the diversity of complex cycles observed globally for pathogens such as M. pneumoniae and parvovirus B19, which may also facilitate co-estimation of environmental sources of stochasticity.
@@
Table 1: Estimated and assumed parameters from fit to 2010-2025 Mycoplasma pneumoniae surveillance data from Denmark.
@@
$\beta_{0}$ Mean transmission rate $0.72/\text{wk}$ $[0.68;0.79]$
$a$ Amplitude of seasonality $0.186$ $[0.174;0.197]$
$\delta$ Rate of waning of immunity $1/(11.2\text{yr})$ $[1/14.4;1/9.8]$
@@
The analogous autonomous system to equations (1)-(3) is obtained by substituting a constant transmission rate $\beta$ in place of $\beta(t)$ (with $\beta$ equal to the mean of its time-varying counterpart). If released away from equilibrium, $I(t)$ performs an underdamped oscillation before the system settles into an equilibrium configurations $(S^{*},I^{*},R^{*})$, where:
$S^{*}=\frac{\gamma+\mu}{\beta},\leavevmode\nobreak\ \leavevmode\nobreak\ I^{*}=\frac{(\beta-\mu-\gamma)(\delta+\mu)}{(\gamma+\delta+\mu)\beta},\leavevmode\nobreak\ \leavevmode\nobreak\ R^{*}=\frac{(\beta-\gamma-\mu)\gamma}{(\gamma+\delta+\mu)\beta}.$
Linearizing the system around this equilibrium and applying a small perturbation, the decay rate and characteristic frequency are given by the eigenvalues of the Jacobian matrix:
@@
We leave out the analytic expressions for the eigenvalues since they are exceedingly long and can be determined from the above matrix by standard methods. With the parameters fitted to the Danish 2010-2025 time series, the eigenvalues are $\lambda_{1}=-2.4\times 10^{-4}$ and $\lambda_{\pm}=-0.0017\pm 0.023i$. Since all real parts are strictly negative, the equilibrium is asymptotically stable.
Bayesian fitting of time series
We fit the seasonal SIRS model as described above using the probabilistic programming language Stan (67, 68), with a sinusoidal seasonally forced transmisson rate $\beta(t)=\beta_{0}\xi(t)(1+a\sin(2\pi t/\text{year}+\phi))$, to positivity rates given in the 2010-2025 data set. $\xi(t)$ is the NPI multiplier which is fixed to 1 until March 2020 and otherwise estimated. The priors used were:
@@
[…]
@@
$r[S\to I]$ =$=\begin{cases}\beta(t)SI/N&\text{if }N>0,\\ 0&\text{otherwise},\end{cases}$
@@
with the time-dependent transmission rate given by $\beta(t)=\beta_{0}(1+a\sin(2\pi t/\text{year}+\phi))$. Denoting the above rates by $r_{i}$, $i\in\{1,\dots,7\}$, the Doob-Gillespie algorithm (50) thus prescribes iteratively propagating the state of the system by drawing a uniformly distributed random number $r_{1}\in[0,1]$, determining the time to next event $\tau=-(\sum_{i}r_{i})^{-1}\log(r_{1})$ and the nature of the event by drawing another random uniform number $r_{2}\in[0,1]$ and determining the smallest $j$ for which $\sum_{i=1}^{j}a_{i}>r_{2}\sum_{i=1}^{7}a_{i}$. Event (or “reaction”) $j$ is then performed and the time is incremented by $\tau$. The procedure is then reiterated until the desired final time is reached.
@@
We thank the Statens Serum Institut (SSI) for providing Danish surveillance data. BFN acknowledges financial support from the Carlsberg Foundation (grant no. CF23-0173). SWP is a Peter and Carmen Lucia Buck Foundation Awardee of the Life Sciences Research Foundation. MHJ is supported by the Novo Nordisk Foundation, grants NNF20OC0064978 and NNF24OC0089788. BTG would like to acknowledge support from Princeton Catalysis and Princeton Precision Health.

### v3

@@ Abstract
The immediate impact of a singular perturbation (a “kick”) to the system (60) is explored in Fig. 4. The perturbation is implemented as either a discrete change in the infected population (Fig. 4A-C) or the susceptible population (Fig. 4D). Since just a single “kick” is given, the system will only transiently settle into a multiannual cycle and we thus measure the strength of the 5-year cycle during the 20 years immediately following the perturbation. In Fig. 4A, we show the trajectory of the system in the $(S,I)$ configuration space immediately before and after a single kick to the system. In this case, the perturbation consists of a sudden reduction in prevalence $I(t)$, but a sudden change in the susceptible population can have a similar effect (Fig. 4D). The annual attractor, where the system resides prior to the perturbation, appears as the central ellipse, while the resulting (transient) five-year cycle encloses it. The rates at which $I(t)$ and $S(t)$ change is reflected in the spacing between the colored markers which are placed one week apart. The colors of the markers themselves indicate the time of year. As is evident, the five-year cycle can be quite complex. The perturbation that initiated the five-year cycle is visible as the approximately vertical line connecting the annual and multiannual cycles, and the state of the system immediately prior to the perturbation is indicated by the red cross. Fig. 4B shows that the timing of the perturbation matters, with a kick during the low-transmission part of the season most likely to excite the multiannual cycle. In other words, a kick of a certain size (whether in the positive or negative direction) has a larger effect if it occurs when circulation is already low. However, Fig. 4C reveals that this seasonal dependence is completely accounted for by the relative magnitude of the perturbation, $\Delta I/I$, i.e. the change in incidence divided by the current incidence at that time. If plotted as a function of this relative magnitude, the strength of the excited cycle shows only a very faint seasonal dependence. Fig. 4D is analogous to 4B, except that the perturbation is applied to the susceptible fraction $S(t)$. Here, seasonal dependence is fairly minimal, since the susceptible fraction fluctuates much less during a season (relative to its mean value) than the prevalence $I(t)$ does.
The mechanism of sustained cycling demonstrated here is fundamentally that of a quasicycle, a mechanism of population cycling proposed in (61). It arises from an otherwise stable, underdamped system being subjected to noise which excites an intrinsic cycle that would otherwise decay. In other words, noise continually disrupts the system’s deterministic approach towards equilibrium. In Supplementary Fig. S8, we demonstrate that the same phenomenon occurs in the absence of seasonal forcing. In the ecological literature, there is a fundamental distinction between phase-remembering and phase-forgetting quasicycles (62, 63, 64). In the former case, the cycle retains phase information over time, such that even temporally well-separated fluctuations exhibit a substantial correlation, typically due to an external periodic forcing. In the phase-forgetting case, this long-term synchronization is eventually lost. In Supplementary Fig. S9, we measure the autocorrelation of the prevalence signal $I(t)$ in the fitted model and find that the multiannual cycles are of the phase-forgetting kind. If one considers only the autocorrelation of the raw prevalence time series (Fig. S9A), one could be led to believe that the system is phase-remembering, since a substantial autocorrelation remains at large lags $\tau$. However, as we show in Fig. S9B, this residual autocorrelation is solely due to the annual modulation, and any phase information in the multiannual cycle is lost over time. We find that the autocorrelation of the multiannual cycle decays approximately exponentially with a half-life of 6-7 years.
The finding that environmental stochasticity can sustain the multiannual cycles of the system is significant because, unlike demographic stochasticity, this mechanism is independent of population sizes.
Aside from explaining the existence of multiannual cycles, the stochastic model also explains the temporary disappearance of the multiannual cycles between the late 1970s and mid-1980s, observed by (2). As shown in the wavelet analysis of model results in Supplementary Fig. S10, the stochastic model exhibits mode-hopping, i.e. spontaneous and temporary reversions to the annual attractor, similar to what can be seen in the wavelet transformed empirical data of Supplementary Fig. S2. The model fit of Fig. 1B also partially captures this phenomenon, with small infrequent changes in the transmission rate in lieu of continuous noise.
@@
Although the deterministic seasonal SIRS model captures the 2010–2025 dynamics well, predicting or forecasting M. pneumoniae incidence may be challenging. We know that long-term predictions using the deterministic model are limited by its tendency to revert to annual dynamics. We have also established that phase information in the multiannual cycles is lost over time when the system is subject to stochasticity, which the real-world system invariably is. Both of these findings suggest that a deterministic model is fundamentally limited in terms of long-term predictive power. In the short term, however, a deterministic model may still be appropriate for forecasting. Fig. 5A shows the trajectory obtained by continuing the deterministic seasonal SIRS model for 10 years beyond the 2010–2025 data (with parameters fitted to said data). However, the low uncertainty of the prediction offered by the deterministic model may be an underestimation. Fig. 5B shows the fit of the seasonal SIRS model with environmental stochasticity. While the 2010–2025 data is still captured well, prediction beyond a year or two is associated with high uncertainty. The noise level was estimated at $\sigma=0.12$ using this model, a degree of stochasticity capable of sustaining pronounced multiannual cycles, as we saw in Fig. 3. We note that this level of stochasticity is substantial, in that $\sigma$ is comparable in magnitude to the estimated amplitude $a$ of seasonality at approx. $0.18$.
The sensitivity to noise can be measured via the Lyapunov exponents of the system. A Lyapunov exponent $\lambda$ measures sensitivity to initial conditions – this also includes the tendency of the system to exponentially amplify ($\lambda>0$) or buffer ($\lambda<0$) perturbations/noise (63). Supplementary Fig. S1 explores the local/instantaneous Lyapunov exponents (LLEs) (65) of the deterministic as well as stochastic model and shows that, while perturbations tend to be transiently amplified, all LLEs eventually become negative, indicating the eventual decay of the perturbation.
While sensitivity to stochastic effects may complicate medium-term predictions (note the spread in predictions in 5B from 2027 onward), some short-term features are relatively robust even at this noise level. Predictions agree that the 2025-2026 season is likely to be a moderate-to-high incidence year, exemplifying how deterministic models yield valuable short-term forecasts even in noise-driven epidemiological systems.
@@
While long-term predictions using the deterministic model are limited by its decay into annual dynamics, the presence of stochasticity limits extended prediction in general, not only by causing variability in the incidence during any given year, but also by enabling mode-hopping, i.e. transitions between annual and multiannual cycles (66, 50), which appear in the empirical data (Supplementary Fig. S2) as well as in model simulations (Supplementary Fig. S10). That the predictions for the years immediately following the large NPI-induced perturbation agree well between the stochastic and deterministic models fits well with the general observation that restorative forces are stronger far from equilibrium, leading the stochastic model to (temporarily) behave more like the deterministic model (46).
@@
In fitting the 2010–2025 time series, we used the PCR positivity rate as an incidence proxy, while we used the number of positive tests for the 1958–1995 data (total number of tests were not on record). Neither of these measures are perfect incidence proxies, and they suffer in distinct ways from biases introduced by varying testing intensities. If testing intensity is increased, this leads to an increase in the number of positive tests, while it leads to a decrease in the test positivity rate (since the denominator – the total number of tests – is increased). It may be possible to enhance fitting further by accounting for fluctuations in testing intensity when such data is available, something we have not attempted to do, beyond simple variance stabilization. Furthermore, the fidelity of the 1958–1995 data series is affected by both the (somewhat imperfect) test characteristics of the cold agglutination and complement fixation tests (since the samples considered positive are those which were positive in both of these tests (2)). We have not taken age structure into account (67), nor have we considered detailed spatial effects which might enhance the impact of demographic stochasticity (68).
@@
The origins of non-trivial cycles in ecological systems and the extent to which stochasticity plays a role in shaping them are long-standing questions in ecology, including epidemiology (60, 69, 70). While the concept of quasicycles – periodic population fluctuations resulting from the interaction of noise with decaying intrinsic cycles – was proposed by (71), the theory has evolved since. Significant analytic advances for SIR-type systems (52) and predator-prey systems (63, 72) were made two decades ago, and quasicycles were proposed as a model of seasonal fluctuations in influenza incidence around the same time (73), albeit without empirical verification. In (53), demographic stochasticity was considered as a potential driver of recurrent outbreaks of avian influenza in North American wild waterfowl and the observed cycles were quantified through wavelet analysis. The authors used a stochastic SIR-type model, extended to include environmental transmission (but not environmental stochasticity), and found that it could explain the occurrence of multiannual outbreak cycles, with the period being largely determined by the intensity of environmental transmission.
The effects of demographic stochasticity have been most widely considered, while environmental noise features less prominently in the epidemiological literature, and studies that consider both are rare. Pertussis (whooping cough) is another example of a disease where multiannual cycles have been studied in detail. In (74), the authors used wavelet analysis to study the multiannual cycles of pertussis across 12 countries and noted that, while a 3–4 year periodicity was typically found (a result echoed by (75)), it is transient in nature. The multiannual cycles of pertussis were previously shown to likely originate from the system’s response to dynamical noise (76), and a study found that stochasticity is essential to even qualitatively capture the cycles of pertussis (60).
For M. pneumoniae, we find that noise-perturbed intrinsic dynamics is sufficient to explain the complex cycles, offering a parsimonious model for the dynamics. Importantly, the model also explains the temporary disappearance of the multiannual cycles observed in the Danish data from the late 1970s until the mid-1980s as a consequence of mode-hopping, i.e. the system reverting to the annual attractor due to a perturbation. This temporary disappearance of the multiannual pattern was previously noted by (2), who proposed changes in daycare attendance as a possible driver.
@@
Earn et al. (77) demonstrated the existence of a rich landscape of stable cycles in the related, seasonally driven SEIR model at high transmission rates (basic reproductive numbers in excess of approx. 4). In the current case of M. pneumoniae however, multiannual cycles are unstable and require stochasticity or perturbations to be sustained. A similar systematic mapping of possible cycles in the seasonal SIRS model – taking stochasticity into account and including subharmonic resonances (71, 78) – would be highly useful.
With the COVID-19 NPIs phased out, another major – if less abrupt – change in M. pneumoniae epidemiology is on the horizon. Macrolide resistance has been growing rapidly in recent years, albeit with large geographical differences (79). It is currently unclear how this will affect epidemiological dynamics, but some studies have found macrolide-resistant strains to be associated with increased virulence and longer disease duration (80, 81), leading us to speculate that average transmissibility could be influenced as well. Extending the current model to incorporate observational data on macrolide resistance would be highly worthwhile, and holds the potential to estimate any effects of macrolide-resistant strains on M. pneumoniae dynamics.
Since the main objective of this study was to determine the mechanism of population cycling in M. pneumoniae, we opted for a fully mechanistic model. This has the further advantage of enabling straight-forward prediction of the overall dynamics in parameter ranges that differ from those obtained from the Danish data. Some parameters are likely intrinsic to the pathogen-host interaction (rates of waning and recovery), while others (transmission rates and birth/death rates) are likely to vary geographically, leading to different dominant periods according to Eq. 5. For adaptive forecasting, however, the formulation of a hybrid semi-mechanistic machine learning model would likely be advantageous. In a recent study (82), a machine learning model was shown to perform well when tasked with short- and medium-term forecasting of historical measles data. The authors note that “[t]he primary focus of mechanistic models has been understanding and characterizing the natural history of transmission […], statistical and machine learning techniques in infectious disease modeling have primarily focused on improving forecasting accuracy without the explicit aim of inferring transmission dynamics.” While beyond the scope of this study, we encourage a synthesis of the two approaches for M. pneumoniae forecasting.
In this study, we have considered both demographic and environmental stochasticity as potential drivers of multiannual outbreak cycles. However, we have not attempted any detailed attribution of stochastic effects to behavioral or demographic sources. A future study could conceivably pinpoint the sources of stochasticity that most strongly shape M. pneumoniae dynamics by extending the current work to incorporate (time-varying) population structure (in terms of age and contact patterns) and mobility data.
In summary, we have offered a parsimonious explanation for the multiannual cycles of M. pneumoniae as a type of quasicycle, and provided a computational and analytic framework for similar analyses of other pathogens and locations. There is a need for comparative studies of the factors governing the diversity of complex cycles observed globally for pathogens such as M. pneumoniae and parvovirus B19, across countries and pathogens. Such comparisons may also enable co-estimation of environmental sources of stochasticity, and the framework presented here is well suited to facilitate this.
@@
Table 1: Estimated and assumed parameters from fit to 2010–2025 Mycoplasma pneumoniae surveillance data from Denmark. parameters labeled “determ.” are estimated using the deterministic seasonal SIRS model, while those labeled “stoch.” are estimated using the seasonal SIRS model with stochastic transmission rate.
@@
$\beta_{0}$ (determ.) Mean transmission rate $0.65/\text{wk}$ $[0.58;0.73]$
$a$ (determ.) Amplitude of seasonality $0.18$ $[0.17;0.19]$
$\delta$ (determ.) Immunity waning rate $(8.2\text{yr}) [1/1$2.0;1/6.4][1/12.0;1/6.4]
$\beta_{0}$ (stoch.) Mean transmission rate $0.71/\text{wk}$ $[0.65;0.77]$
$a$ (stoch.) Amplitude of seasonality $0.18$ $[0.15;0.20]$
$\delta$ (stoch.) Immunity waning rate $(8.7\text{yr}) [1/1$2.0;1/6.9][1/12.0;1/6.9]
$\sigma$ (stoch.) Level of stochasticity $0.12$ $[0.11;0.13]$
@@
The analogous autonomous system to Eqs. 2-4 is obtained by substituting a constant transmission rate $\beta$ in place of $\beta(t)$. If perturbed away from equilibrium, $I(t)$ performs an underdamped oscillation while approaching the equilibrium configuration $(S^{*},I^{*},R^{*})$, where:
$S^{*}=\frac{\gamma+\mu}{\beta},\penalty 10000\ \penalty 10000\ I^{*}=\frac{(\beta-\mu-\gamma)(\delta+\mu)}{(\gamma+\delta+\mu)\beta},\penalty 10000\ \penalty 10000\ R^{*}=\frac{(\beta-\gamma-\mu)\gamma}{(\gamma+\delta+\mu)\beta}.$
Linearizing the system around this equilibrium and applying a small perturbation, the decay rate and characteristic frequency are given by the real and imaginary parts of the eigenvalues of the Jacobian matrix:
@@
The angular frequency of the intrinsic oscillation is then given by $\omega=\sqrt{\Omega_{0}^{2}-(\Gamma/2)^{2}}$, where $\Omega_{0}^{2}=(\beta-\gamma-\mu)(\delta+\mu)$ and $\Gamma=\frac{(\beta+\delta)(\delta+\mu)}{\gamma+\delta+\mu}$. Since $\beta-\gamma$ is much greater than $\mu$ and $\delta$, it holds that $\omega\approx\Omega_{0}$, although we choose to keep $\Gamma$ in our calculations. The corresponding period is then given by
$T=\frac{2\pi}{\sqrt{\Omega_{0}^{2}-(\Gamma/2)^{2}}}.$ (5)
With the parameters fitted to the Danish 2010–2025 time series, the eigenvalues are $\lambda_{1}=-2.4\times 10^{-4}\text{wk}^{-1}$ and $\lambda_{\pm}=(-0.0020\pm 0.024i)\text{wk}^{-1}$. Since all real parts are strictly negative, the equilibrium is asymptotically stable.
Bayesian fitting of a deterministic model to 2010–2025 data
We fit the seasonal SIRS model as described above using the probabilistic programming language Stan (83, 84), with a sinusoidal seasonally forced transmission rate $\beta(t)=\beta_{0}\xi(t)(1+a\sin(2\pi t/\text{year}+\phi))$, to positivity rates given in the 2010–2025 data set. $\xi(t)$ is the NPI multiplier which is fixed to 1 until March 2020 and otherwise estimated. The priors used were:
@@
Here $x_{\text{data}}$ and $x_{\text{model}}$ refer to the measured incidence proxy (positivity rate, arbitrarily rescaled) and the modeled incidence, respectively. Note that for the parameters $\beta_{0}$, $\rho$, $S(t=0)$, $a$, $\delta$, $x_{\text{data}}$, $\sigma_{\text{obs}}$ and the NPI multipliers, only the part of the normal distribution with positive support is used, since the parameters in question are necessarily non-negative. The fit to the 1958–1995 MpCF-based data follows similar principles and is detailed in the appendix. For the purposes of fitting, the model (Eqs. 2-4) is implemented using the efficient discretization scheme described in (25), with a time-step of $0.125\text{wk}$. This discretization was used for Figs. 1, 3, 4, 5, S4, S5, S7, S8 and S10, whereas the calculations of Figs. S1 and S9 were performed using standard Euler integration of Eqs. 2-4. Convergence was assessed based on the absence of diagnostic warnings from CmdStanPy (84), including the absence of divergent chains, no exceedances of tree-depth, sufficient E-BFMI (>0.3>0.3), and satisfactory $\hat{R}$ (<1.05<1.05) and effective sample size (bulk $\text{ESS}>100\times\text{\#chains}$) for all parameters.
Bayesian fitting of a stochastic model
We fit a seasonal SIRS model with environmental stochasticity in the form of Gamma-distributed multiplicative noise in the transmission rate, as used in Fig. 5. The model set-up, using Stan (83), closely follows that of the deterministic model. A new fitted parameter, the noise level $\sigma$, is introduced which controls the magnitude of noise in $\beta(t)$, with prior $\sigma\sim N(0.1,0.02)$ (positive-support part of normal distribution only). The individual perturbations are then implemented by letting $\beta(t)\to\eta(t)\beta(t)$ where $\eta\sim\Gamma[1/\sigma^{2},\sigma^{2}]$ with the notation $\Gamma[\alpha,\theta]$ referring to a Gamma distribution with shape $\alpha$ and scale $\theta$. For computational efficiency, we only draw a new value of $\eta(t)$ every other week.
@@
$r[S\to I]$ $=\begin{cases}\beta(t)SI/N&\text{if }N>0,\\ 0&\text{otherwise},\end{cases}$
@@
with the time-dependent transmission rate given by $\beta(t)=\beta_{0}(1+a\sin(2\pi t/\text{year}+\phi))$. Note that here, $S$, II, and $R$ denote the populations of susceptible, infected, and recovered individuals (as integers), not the population fractions. Denoting the above rates by $r_{i}$, $i\in\{1,\dots,7\}$, the Doob-Gillespie algorithm (54) thus prescribes iteratively propagating the state of the system by drawing a uniformly distributed random number $u_{1}\in[0,1]$, determining the time to next event $\tau=-(\sum_{i}r_{i})^{-1}\log(u_{1})$ and the nature of the event by drawing another random uniform number $u_{2}\in[0,1]$ and determining the smallest $j$ for which $\sum_{i=1}^{j}r_{i}>u_{2}\sum_{i=1}^{7}r_{i}$. Event (or “reaction”) $j$ is then performed and the time is incremented by $\tau$. The procedure is then reiterated until the desired final time is reached. The metapopulation simulations of Supplementary Fig. S11 use the same algorithm, but with modified rates. Details are given in the supplement.
@@
We thank Statens Serum Institut (SSI) for providing Danish surveillance data. We would also like to thank Mercedes Pascual for highly valuable input. BFN acknowledges financial support from the Carlsberg Foundation (grants CF23-0173 and CF24-1337). The PandemiX Center is funded by the Danish National Research Foundation (grant DNRF170). SWP is a Peter and Carmen Lucia Buck Foundation Awardee of the Life Sciences Research Foundation. MHJ is supported by the Novo Nordisk Foundation, grants NNF20OC0064978 and NNF24OC0089788. BTG would like to acknowledge support from Princeton Catalysis Initiative and Princeton Precision Health.